# Approaching Constrained Markov Decision Process

### v1

@@ 3.1 Overview of Our Approach
For our approach, we identify the optimal actions and then stick to the optimal actions. With the linear programming reformulations (8), identifying the optimal actions corresponds to finding an optimal solution of the linear program. However, there can be multiple optimal solutions of the LP (8). We characterize one optimal solution and we exploit its structure during the entire horizon. We first observe the following result regarding an optimal solution to the LP (8), which follows from standard LP theory. The proof of Lemma 3.1 is relegated to Section 8.
@@
Denote by $b$ the number of rows in the matrix $B$. Then, there exists a subset $\mathcal{J}\subset[K]$ with $K^{\prime}=|\mathcal{J}|$ and an optimal solution $\bm{q}^{*}$ to LP (8) such that there are $b+K^{\prime}$ variables in $\bm{q}^{*}$ are non-zero. Moreover, we denote by $\mathcal{I}$ the index set of the non-zero element in $\bm{q}^{*}$. Then, the optimal solution $\bm{q}^{*}$ can be uniquely determined as the solution to the linear system
$C(\mathcal{J},\mathcal{I})\bm{q}_{\mathcal{I}}=\bm{\alpha}_{\mathcal{J}},$ (9a)
$B(:,\mathcal{I})\bm{q}_{\mathcal{I}}=\bm{\mu},$ (9b)
$\bm{q}_{\mathcal{I}^{c}}=0.$ (9c)
with $\mathcal{I}^{c}$ being the complementary set of the index set $\mathcal{I}$.
In the first part, we aim to identify the set $\mathcal{I}$ and $\mathcal{J}$. By doing so, we also obtain the optimal actions to take given each state (and each period), by noting the relationship between the optimal actions and the optimal solution (occupancy measure) in (7). In the second part, we aim to obtain the optimal solution (occupancy measure) by solving the linear equations (3.1). However, since the parameters of the linear equations (3.1) are unknown, we cannot directly solve (3.1). Instead, we need to use historical samples to construct estimates of the parameters of (3.1) and gradually approximate the optimal solution. To this end, we seek tools from the online packing/linear programming literature (e.g. Agrawal et al. (2014), Li and Ye (2022), Vera and Banerjee (2021), Bumpensanti and Wang (2020)), where we interpret the right-hand side (RHS) of the linear equations (3.1) as resource capacities. Note that a crucial element in deriving the logarithmic or better problem-dependent regret in the previous literature is to let our algorithm be adaptive to the remaining resource capacities in real time. We adopt this adaptive resolving idea in our algorithm.
In the remaining part, we first describe how to construct estimates of unknown parameters for the CMDP problem and we present our formal algorithms. We show how to identify one optimal basis, and how to approximate the optimal solution with an adaptive constraint design.
@@ 4 Construct Estimates in CMDP
with $\lambda_{N_{0}}$ being a parameter that we specify later for both the infinite horizon discounted problem and the finite horizon episodic problem. To bound the estimation gap, it is useful to bound the optimal dual solution to (8). To this end, we adopt the approach in Hiriart-Urruty and Lemaréchal (1996), Nedić and Ozdaglar (2009) that utilizes Slater’s condition, which is imposed as an assumption below. {assumption} There exists a policy $\bar{\pi}$ such that all the resource constraints are satisfied strictly. In other words, there exists an occupancy measure $\bar{\bm{q}}$ such that $B\bar{\bm{q}}=\bm{\mu}$ and $C\bar{\bm{q}}<\bm{\alpha}$. In fact, for each state $s\in\mathcal{S}$, there exists a null action that consumes no resource. The Slater point $\bar{\bm{q}}$ can be set as the policy that takes the null action given each state. The estimation error will be related to the gap between the Slater point $\bar{\bm{q}}$ and the optimal point $\bm{q}^{*}$. We then define the lower gap as
@@ 4.1 Estimates for the Infinite Horizon Discounted Problem
Following the standard Hoeffding’s inequality, we know that $|\bar{r}_{N_{0}}(s,a)-\hat{r}(s,a)|$, $|\bar{c}_{k,N_{0}}(s,a)-\hat{c}_{k}(s,a)|$, and $|\bar{P}_{N_{0}}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|$ are all upper bounded by $\mathrm{Rad}(N_{0},\varepsilon)$ with probability at least $1-\varepsilon$. Then, we use the following LP with empirical estimates to approximate the value of $V^{\mathrm{Infi}}$, as given in (10). Here, $\bar{C}_{N_{0}}\in\mathbb{R}^{K\times|\mathcal{S}|}$, and the element at the $k$-th row and $(s,a)$-th column of the matrix $\bar{C}_{N_{0}}$ is $\bar{c}_{k,N_{0}}(s,a)$. $\bar{B}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$, and the element at $s$-th row and $(s^{\prime},a^{\prime})$-th column of the matrix $\bar{B}_{N_{0}}$ is $(\delta_{s,s^{\prime}}-\gamma\cdot\bar{P}_{N_{0}}(s|s^{\prime},a^{\prime}))$. $\bar{\bm{r}}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ with the $(s,a)$-th element being $\bar{r}_{N_{0}}(s,a)$. We have the following result, where the proof is relegated to Section 9.
@@ 5.1 Characterize One Optimal Basis
3:Initialize $\mathcal{I}$ to be the whole index set that contains every row index of matrix $B$ in (8) and $\mathcal{J}=[K]$.
@@
11:     Compute the value of $\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}$ as in (5.1).
@@
where $\mathcal{I}^{c}$ denotes the complementary set of $\mathcal{I}$. Note that if $V-V_{\mathcal{I}}>0$, we know that $\mathcal{I}^{c}$ contains a non-zero basic variable. The steps 4-8 in Algorithm 1 reflect this point. Starting from $\mathcal{I}$ denoting the whole index set, we sequentially delete one element $i$ (denoting $(s,a)$ in the infinite horizon discounted problem and $(s,a,h)$ in the finite horizon episodic problem) from the set $\mathcal{I}$. Once we detected that $V-V_{\mathcal{I}\backslash\{i\}}>0$, we know that $i$ is a non-zero basic variable and we add $i$ back into the set $\mathcal{I}$. In this way, we can classify all the basic variables into the set $\mathcal{I}$. However, since we do not know the exact value of $V$ and $V_{\mathcal{I}}$, we use the estimates and compare the value of $\bar{V}_{N_{0}}$ and $\bar{V}_{\mathcal{I},N_{0}}$. For this comparison to be valid, the estimation error has to be smaller than the intrinsic gap between $V$ and $V_{\mathcal{I}}$. We define a constant
@@ 5.2 Our Formal Algorithm
Remark. Note that the adaptive resolving step (23) is akin to the idea of resolving an LP in online LP, Bandits with Knapsacks, online allocation problems, where logarithmic or better regret is obtained in a series of work Flajolet and Jaillet (2015), Arlotto and Gurvich (2019), Li and Ye (2022), Li et al. (2021b), Vera and Banerjee (2021), Vera et al. (2021), Bumpensanti and Wang (2020), Jiang et al. (2022). However, since we have identified the optimal basis $\mathcal{I}$ and $\mathcal{J}$ and we stick to this optimal basis to avoid the non-degeneracy assumption, resolving the LP would become resolving the linear equations in (23).
@@ 6.1 High Probability Bound for Identifying One Optimal Basis
We then denote by $\{\sigma_{1}(A^{*}),\dots,\sigma_{|\mathcal{S}|+K^{\prime}}(A^{*})\}$ the eigenvalues of the matrix $A^{*}$. We define $\sigma$ as
@@
From the non-singularity of the matrix $A$, we know that $\sigma>0$.
Remark. It is clear to see that $\sigma$ in (28) relates to the conditional number of the matrix $A^{*}$. In fact, it holds that $\kappa(A^{*})\leq\frac{|\mathcal{S}|+K}{\sigma}$, where $\kappa(A^{*})$ denotes the conditional number of the matrix $A^{*}$. The value of $\sigma$ will show up in our final bounds as it characterizes how hard it is to learn the matrix $A^{*}$. Also, $\sigma$ gives a natural upper bound of the range of the optimal dual variable corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. Note that in previous literature that establishes logarithmic regret for online LP/resource allocation (e.g. Vera and Banerjee (2021), Li et al. (2021b), Jiang et al. (2022)), the conditional number of the constraint matrix will also show up in the final bounds. Therefore, we regard the existence of $\sigma$ as the consequence of adopting the online LP/resource allocation framework. On the other hand, for the MDP problem without resource constraints, the matrix $A^{*}$ can simply be represented as the probability transition matrix and one can easily show that $\sigma\geq 1-\gamma$. We will use this bound when deriving our guarantees for MDP problems without resource constraints.
@@ 6.2 Regret Bounds
$N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\delta^{2}\}}\cdot\frac{\log(N)}{N}\right)$
where the parameters $\alpha=\min_{k\in[K]}\{\alpha_{k}\}$, $\delta=\min\{\delta_{1},\delta_{2}\}$ with $\delta_{1}$ given in (18) and $\delta_{2}$ given in (21), $\sigma$ given in (28). Also, for any $k\in[K]$, we have
$N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\mathbb{E}\left[\bm{q}^{n}\right]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\delta^{2}\}}\cdot\frac{\log(N)}{N}\right).$ (33)
@@ 6.3 Sample Complexity Bounds
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\delta^{2}\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
@@
$O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\xi\cdot\delta^{2}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$ (39)
@@
Note that one significant term in our sample complexity bound in Proposition 6.5 is $\xi\cdot\delta^{2}$. Recall that $\xi$ represents the minimal mass of the optimal occupancy measure and $\delta$ represents the change of the objective value of LP (8) if we restrict one non-zero basic variable in an optimal solution to be 00. In this sense, $\xi\cdot\delta^{2}$ can be understood as the gap between the value of the optimal policy and the value of a sub-optimal policy that is restricted from taking the optimal action given a state. Such a problem-dependent gap shows up broadly in the literature for online decision-making with (logarithmic) problem-dependent guarantees, for example, in the multi-arm-bandit problem (Lai and Robbins 1985) and the reinforcement learning problem (Auer et al. 2008).

### v2

@@ 3.1 Characterization of Instance Hardness
Note that one crucial step for achieving problem-dependent bounds is to characterize the hardness of the underlying problem instance and define a gap that separates the optimal policies from the others. For multi-arm-bandit problem, the characterization of the hardness can be the gap between the optimal arm and the best sub-optimal arm (e.g. Lai and Robbins (1985)). For reinforcement learning problem, the characterization of the hardness can also be the gap between the optimal policy and the best sub-optimal policy (e.g. Auer et al. (2008)). As long as this separation gap is a positive constant, denoted by $\Delta$, separating the optimal policy from the others with a probability at least $1-\epsilon$ would require samples at most $\frac{1}{\Delta}\cdot\log(1/\epsilon)$, which finally implies the instance-optimal sample complexity bound.
For CMDP problem, characterizing the hardness of the problem instance can be hard. Based on LP (8), we know that a feasible policy corresponds to a feasible solution and the sub-optimal solution can be arbitrarily close to the optimal solution since the feasible set is “continuous”. Therefore, there is no direct way to identify a positive gap between the optimal policies and the sub-optimal ones. However, from standard LP theory, we know that one corner point of the feasible region must be one optimal solution. Therefore, we can simply focus on the corner points when solving LP (8) and we define the gap as the distance between the optimal corner point and the sub-optimal corner point, as illustrated in Figure 1. As we will show later, the LP reformulation (8) and such a characterization of hardness via corner points will inspire our entire approach.
Figure 1: A graph illustration of the hardness characterization via LP basis, where the shaded area denotes the feasible region for the policies.
Since the problem hardness is characterized via corner points, it is essential to provide further characterization of the corner points. Note that in LP theory, the corner point is called basic solution and can be represented by LP basis, which involves the set of basic variables that are allowed to be non-zero, and the set of active constraints that are binding under the corresponding basic solution. For our approach, we identify one optimal basic solution and then stick to this optimal basic solution. However, there can be multiple optimal bases of the LP (8). We characterize one optimal basis and we exploit its structure during the entire horizon. We first observe the following result regarding an optimal basis to the LP (8), which follows from standard LP theory. The proof of Lemma 3.1 is relegated to Section 9 for completeness.
@@
Denote by $b$ the number of rows in the matrix $B$. Then, there exists a subset $\mathcal{J}^{*}\subset[K]$ with $K^{\prime}=|\mathcal{J}^{*}|$ and an optimal solution $\bm{q}^{*}$ to LP (8) such that there are $b+K^{\prime}$ variables in $\bm{q}^{*}$ are non-zero. Moreover, we denote by $\mathcal{I}^{*}$ the index set of the non-zero element in $\bm{q}^{*}$. Then, the optimal solution $\bm{q}^{*}$ can be uniquely determined as the solution to the linear system
$C(\mathcal{J}^{*},\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\alpha}_{\mathcal{J}^{*}},$ (9a)
$B(:,\mathcal{I}^{*})\bm{q}_{\mathcal{I}^{*}}=\bm{\mu},$ (9b)
$\bm{q}_{\mathcal{I}^{*c}}=0.$ (9c)
with $\mathcal{I}^{*c}$ being the complementary set of the index set $\mathcal{I}^{*}$.
3.2 Overview of Our Approach
In the first part, we aim to identify the set $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$, which characterizes one optimal basis of LP (6). By doing so, we also obtain the optimal actions to take given each state (and each period), by noting the relationship between the optimal actions and the optimal solution (occupancy measure) in (7). In the second part, we aim to obtain the optimal solution (occupancy measure) by solving the linear equations (9). However, since the parameters of the linear equations (9) are unknown, we cannot directly solve (9). Instead, we need to use historical samples to construct estimates of the parameters of (9) and gradually approximate the optimal solution. To this end, we seek tools from the online packing/linear programming literature (e.g. Agrawal et al. (2014), Li and Ye (2022), Vera and Banerjee (2021), Bumpensanti and Wang (2020)), where we interpret the right-hand side (RHS) of the linear equations (9) as resource capacities. Note that a crucial element in deriving the logarithmic or better problem-dependent regret in the previous literature is to let our algorithm be adaptive to the remaining resource capacities in real-time. We adopt this adaptive resolving idea in our algorithm.
In the remaining sections, we first describe how our LP-based approach implies a characterization of the “hardness” of the underlying problem instance. We then describe how to construct estimates of unknown parameters for the CMDP problem and we present our formal algorithms. We show how to identify one optimal basis, and how to approximate the optimal solution with an adaptive constraint design.
@@ 4 Construct Estimates in CMDP
with $\lambda_{N_{0}}$ being a parameter that we specify later. To bound the estimation gap, it is useful to bound the optimal dual solution to (8). To this end, we adopt the approach in Hiriart-Urruty and Lemaréchal (1996), Nedić and Ozdaglar (2009) that utilizes Slater’s condition, which is imposed as an assumption below. {assumption} There exists a policy $\bar{\pi}$ such that all the resource constraints are satisfied strictly. In other words, there exists an occupancy measure $\bar{\bm{q}}$ such that $B\bar{\bm{q}}=\bm{\mu}$ and $C\bar{\bm{q}}<\bm{\alpha}$. In fact, for each state $s\in\mathcal{S}$, there exists a null action that consumes no resource. The Slater point $\bar{\bm{q}}$ can be set as the policy that takes the null action given each state. The estimation error will be related to the gap between the Slater point $\bar{\bm{q}}$ and the optimal point $\bm{q}^{*}$. We then define the lower gap as
@@ 4.1 Estimates for the Infinite Horizon Discounted Problem
Following the standard Hoeffding’s inequality, we know that $|\bar{r}_{N_{0}}(s,a)-\hat{r}(s,a)|$, $|\bar{c}_{k,N_{0}}(s,a)-\hat{c}_{k}(s,a)|$, and $|\bar{P}_{N_{0}}(s^{\prime}|s,a)-P(s^{\prime}|s,a)|$ are all upper bounded by $\mathrm{Rad}(N_{0},\varepsilon)$ with probability at least $1-\varepsilon$. Then, we use the following LP with empirical estimates to approximate the value of $V^{\mathrm{Infi}}$, as given in (10). Here, $\bar{C}_{N_{0}}\in\mathbb{R}^{K\times|\mathcal{S}||\mathcal{A}|}$, and the element at the $k$-th row and $(s,a)$-th column of the matrix $\bar{C}_{N_{0}}$ is $\bar{c}_{k,N_{0}}(s,a)$. $\bar{B}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}|\times|\mathcal{S}||\mathcal{A}|}$, and the element at $s$-th row and $(s^{\prime},a^{\prime})$-th column of the matrix $\bar{B}_{N_{0}}$ is $(\delta_{s,s^{\prime}}-\gamma\cdot\bar{P}_{N_{0}}(s|s^{\prime},a^{\prime}))$. $\bar{\bm{r}}_{N_{0}}\in\mathbb{R}^{|\mathcal{S}||\mathcal{A}|}$ with the $(s,a)$-th element being $\bar{r}_{N_{0}}(s,a)$. We have the following result, where the proof is relegated to Section 10.
@@ 5.1 Characterize One Optimal Basis
3:Initialize $\mathcal{I}$ to be the whole index set that contains every column index of matrix $B$ in (8) and $\mathcal{J}=[K]$.
@@
11:     Compute the value of $\text{Dual}_{\mathcal{J}^{\prime},\mathcal{I},N_{0}}$ as in (20).
@@
where $\mathcal{I}^{c}$ denotes the complementary set of $\mathcal{I}$. Note that if $V-V_{\mathcal{I}}>0$, we know that $\mathcal{I}^{c}$ contains a non-zero basic variable. The steps 4-8 in Algorithm 1 reflect this point. Starting from $\mathcal{I}$ denoting the whole index set, we sequentially delete one element $i$ (denoting $(s,a)$ in the infinite horizon discounted problem) from the set $\mathcal{I}$. Once we detected that $V-V_{\mathcal{I}\backslash\{i\}}>0$, we know that $i$ is a non-zero basic variable and we add $i$ back into the set $\mathcal{I}$. In this way, we can classify all the basic variables into the set $\mathcal{I}$. However, since we do not know the exact value of $V$ and $V_{\mathcal{I}}$, we use the estimates and compare the value of $\bar{V}_{N_{0}}$ and $\bar{V}_{\mathcal{I},N_{0}}$. For this comparison to be valid, the estimation error has to be smaller than the intrinsic gap between $V$ and $V_{\mathcal{I}}$. We define a constant
@@ 5.2 Our Formal Algorithm
Remark. Note that the adaptive resolving step (23) is akin to the idea of resolving an LP in online LP, Bandits with Knapsacks, online allocation problems, where logarithmic or better regret is obtained in a series of work Flajolet and Jaillet (2015), Arlotto and Gurvich (2019), Li and Ye (2022), Li et al. (2021c), Vera and Banerjee (2021), Vera et al. (2021), Bumpensanti and Wang (2020), Jiang et al. (2022). However, since we have identified the optimal basis $\mathcal{I}$ and $\mathcal{J}$ and we stick to this optimal basis to avoid the non-degeneracy assumption, resolving the LP would become resolving the linear equations in (23).
@@ 6.1 High Probability Bound for Identifying One Optimal Basis
We then denote by $\{\sigma_{1}(A^{*}),\dots,\sigma_{|\mathcal{S}|+K^{\prime}}(A^{*})\}$ the eigenvalues of a matrix $A^{*}$. We define $\sigma$ as
@@
From the non-singularity of the matrix $A^{*}$, we know that $\sigma>0$.
Remark. It is clear to see that $\sigma$ in (28) relates to the conditional number of the matrix $A^{*}$. In fact, it holds that $\kappa(A^{*})\leq\frac{|\mathcal{S}|+K}{\sigma}$, where $\kappa(A^{*})$ denotes the conditional number of the matrix $A^{*}$. The value of $\sigma$ will show up in our final bounds as it characterizes how hard it is to learn the matrix $A^{*}$. Also, $\sigma$ gives a natural upper bound of the range of the optimal dual variable corresponding to the optimal basis $\mathcal{I}^{*}$ and $\mathcal{J}^{*}$. Note that in previous literature that establishes logarithmic regret for online LP/resource allocation (e.g. Vera and Banerjee (2021), Li et al. (2021c), Jiang et al. (2022)), the conditional number of the constraint matrix will also show up in the final bounds. Therefore, we regard the existence of $\sigma$ as the consequence of adopting the online LP/resource allocation framework. On the other hand, for the MDP problem without resource constraints, the matrix $A^{*}$ can simply be represented as the probability transition matrix and one can easily show that $\sigma\geq 1-\gamma$. We will use this bound when deriving our guarantees for MDP problems without resource constraints.
@@ 6.2 Regret Bounds
$N\cdot V_{r}(\pi^{*},\mu_{1})-\sum_{n=1}^{N}\hat{\bm{r}}^{\top}\mathbb{E}[\bm{q}^{n}]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log(N)}{N}\right)$
where the parameters $\alpha=\min_{k\in[K]}\{\alpha_{k}\}$, $\Delta=\min\{\delta^{2}_{1},\delta^{2}_{2}\}$ with $\delta_{1}$ given in (18) and $\delta_{2}$ given in (21), $\sigma$ given in (28). Also, for any $k\in[K]$, we have
$N\cdot\alpha_{k}-\sum_{n=1}^{N}\hat{\bm{c}}_{k}^{\top}\mathbb{E}\left[\bm{q}^{n}\right]\leq O\left(\frac{(|\mathcal{S}|+K)^{3}}{\alpha\cdot\sigma\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log(N)}{N}\right).$ (33)
@@ 6.3 Sample Complexity Bounds
$O\left(\frac{(|\mathcal{S}|+K)^{3}\cdot|\mathcal{S}|\cdot|\mathcal{A}|}{\alpha^{2}\cdot\xi\cdot\sigma(1-\gamma)\cdot\min\{\sigma^{2},(1-\gamma)^{2}\cdot\Delta\}}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$
@@
$O\left(\frac{|\mathcal{S}|^{4}\cdot|\mathcal{A}|}{(1-\gamma)^{4}\cdot\xi\cdot\Delta}\cdot\frac{\log^{2}(1/\varepsilon)}{\varepsilon}\right),$ (39)
@@
Note that one significant term in our sample complexity bound in Proposition 6.5 is $\xi\cdot\Delta$. Recall that $\xi$ represents the minimal mass of the optimal occupancy measure and $\delta$ represents the change of the objective value of LP (8) if we restrict one non-zero basic variable in an optimal solution to be 00. In this sense, $\xi\cdot\Delta$ can be understood as the gap between the value of the optimal policy and the value of a sub-optimal policy that is restricted from taking the optimal action given a state, corresponding to the instance hardness via corner point as described in Section 3.1. Such a problem-dependent gap shows up broadly in the literature for online decision-making with (logarithmic) problem-dependent guarantees, for example, in the multi-arm-bandit problem (Lai and Robbins 1985) and the reinforcement learning problem (Auer et al. 2008).